# relating real and p-adic Kazhdan-Lusztig polynomials

### v1

@@ Abstract.
Fix an integral semisimple element $\lambda$ in the Lie algebra $\mathfrak{g}$ of a complex reductive algebraic group $G$. Let $L$ denote the centralizer of $\lambda$ in $G$ and let $\mathfrak{g}(-1)$ denote the $-1$-eigenspace of $\mathop{\hbox{ad}}\nolimits(\lambda)$ in $\mathfrak{g}$. Under a natural hypothesis (which is always satisfied for classical subgroups of $\mathrm{GL}(n)$), we embed the closure of each $L$ orbit on $\mathfrak{g}(-1)$ into the closure of an orbit of a symmetric subgroup $K$ containing $L$ on a partial flag variety for $G$. We use this to relate the local intersection homology of the later orbit closures to the former orbit closures. This, in turn, relates multiplicity matrices for split real and $p$-adic groups. We also describe relationships between “microlocal packets” of representations of these groups.
@@
On the other hand, let $K$ denote the identity component of the fixed points in $G$ of the automorphism $\theta$ obtained by conjugation by $\exp(i\pi\lambda)$. Then $K$ acts with finitely many orbits on the partial flag variety $\mathcal{P}$ consisting of conjugates of the sum of the nonnegative eigenspaces of $\mathop{\hbox{ad}}\nolimits(\lambda)$. Once again, one can consider local intersection homology Poincaré polynomials. Vogan [Vo3] described a finite effective algorithm to compute them which has been implemented in great generality in the software package atlas. Evaluating these polynomials at 1 gives multiplicities of irreducibles in standard representation of the split real form of the Langlands dual of $G$ [Vo4, ABV].
@@
Equation (1.1) generalizes the main geometric result for $\mathrm{GL}(n)$ of Ciubotaru-Trapa [CT]; see Remark 3.3. In Remark 3.4, we also give analogous results equating the $p$-adic polynomials with Kazhdan-Lusztig polynomials arising from category $\mathcal{O}$, generalizing results of Zelevinsky [Z] for $\mathfrak{g}\mathfrak{l}(n)$ to all classical groups.
@@
Note also that we have assumed $K$ to be the identity component of the fixed points of $\theta$. This is perhaps unnatural from the point of view of representations of algebraic groups, and requires explanation. If we had instead worked with the potentially disconnected $K^{\prime}=G^{\theta}$, the orbits of $K^{\prime}$ on $\mathcal{P}$ can be reducible, and different irreducible components can contribute to the local intersection cohomology in ways that make the right-hand side of (1.1) a summand of the left-hand side. In any particular case, this is tractable to understand, but general statements in the presence of this kind of disconnectedness are somewhat cumbersome. See Remark 3.5.
@@
Let $G=\mathrm{GL}(n,{\mathbb{C}})$ and let $\mathfrak{t}$ denote the diagonal Cartan subalgebra. After a central shift, the integrality of $\lambda$ implies that we may assume $\lambda$ consists of integer entries. Then $K\simeq\mathrm{GL}(p,\mathbb{C})\times\mathrm{GL}(q,\mathbb{C})$ where $p$ is the number of even entries of $\lambda$ and $q$ is the number of odd entries. The symmetric pair $(G,K)$ corresponds to the real group $\mathrm{U}(p,q)$. This is the setting of [CT].
@@
$\mathfrak{g}(-1)=\bigoplus_{i=1}^{\ell}\left[\mathfrak{l}_{i}\cap\overline{\mathfrak{u}}{}\right]$ (2.2)
@@
This is the map that is used in [CT] to define $\mathcal{O}\mapsto Q_{\mathcal{O}}$. But note that we may take any ordering of the set $\mathscr{P}$ and thus define different maps $\mathcal{O}\mapsto Q_{\mathcal{O}}$ which will have the properties that we describe in Theorem 3.1. ∎
@@
Let $G$ be simple of type F4. Suppose $\lambda$ is one-half the middle element of a Jacobson-Morozov triple for the nilpotent orbit labeled in the Bala-Carter classification by F4(a3). The weighted Dynkin diagram of this orbit (in the standard Bourbaki order) is 0200. Thus $\mathfrak{p}{}$ is the maximal parabolic corresponding the long middle root on the Dynkin diagram of F4. The opposite nilradical $\overline{\mathfrak{u}}{}$ is a three-step nilpotent algebra whose first step is $\mathfrak{g}(-1)$. Since $\mathfrak{p}{}$ is maximal, the only possibility is for $\mathscr{P}$ to consist of a single algebra, namely all of $\mathfrak{g}$. But we have already remarked that $\mathfrak{g}\cap\overline{\mathfrak{u}}{}=\bar{\mathfrak{u}}{}$ properly contains $\mathfrak{g}(-1)$. Thus there is no choice of $\mathscr{P}$ that satisfies (2.2)∎
@@ 3.2. Statement of Main Results
[…]
@@
When $G=\mathrm{GL}(n,\mathbb{C})$ and the choice of $\mathscr{P}$ (and its ordering) is the one described in Example 2.5, then (3.6) in Theorem 3.1 is [CT, Theorem 2.5]. ∎
@@
Note, however, that there are no nontrivial local systems in $\Xi(P,Y)$. So there is no hope in matching polynomials that arise for nontrivial local systems for $L$ orbits on $\mathfrak{g}(-1)$ with classical Kazdan-Lusztig polynomials; cf. Remark 3.2. This is a reason for using $K$ orbits in order to match more general polynomials for $\mathfrak{g}(-1)$. ∎
@@
and that the analogous conclusion of (3.7) also holds. If we do not assume $Q_{\gamma}$ is irreducible, then (3.9) can fail; see [BT, Remark 3.6] for a discussion of the an example in $\mathrm{SO}(8)$. The one place in the argument below that we need $Q_{\gamma}$ to be irreducible is to deduce the conclusion of Corollary 4.4.
@@ 4.3. Induced bundles
[…]
@@
In this section, we study a special class of examples that includes some cases (like Example 2.7) where there does not exist a set $\mathscr{P}$ satisfying (2.1) and (2.2). In these cases, we do not need to truncate the exponential map. Alternatively, one can think of this section as a kind of basic case, and the definition of $\epsilon$ in Definition 2.2 as an iteration of these cases using the bundle constructed in (4.5). This can be made precise, but isn’t necessary for our purposes here.
@@
[…]
@@
If $K$ is connected, $[K\cap\bar{B}]\cdot\epsilon^{\prime}(x)$ is open and dense in $K\cdot\epsilon^{\prime}(x)$. In general, $[K\cap\bar{B}]\cdot\epsilon^{\prime}(x)$ is open and dense is a subset of the irreducible components of $K\cdot\epsilon^{\prime}(x)$.
@@
To conclude, we sketch some consequences related to micro-packets of representations. Part of our motivation is to explain the appearance of the Kashiwara-Saito singularity discovered by Cunningham-Fiora-Kitt [CuFK].
@@
Meanwhile, if we assume further that $\lambda$ is hyperbolic, then according to Lusztig’s classification of unipotent representation of graded affine Hecke algebras ([Lu1, Lu2, Lu3]), the set $\Xi(L,\mathfrak{g}(-1))$ parameterizes certain unipotent representations of the split $F=\mathbb{Q}_{p}$ form $G^{\prime}_{F}$ of the Langlands dual group $G^{\vee}$. (The subset $\Xi_{0}(L,\mathfrak{g}(-1))$ consisting of irreducible local systems of Springer type parametrizes irreducible unramified representations of $G^{\prime}_{F}$.) Write $\pi_{F}(\gamma)$ for the irreducible unipotent representation of $G^{\prime}_{F}$ corresponding to $\gamma\in\Xi(L,\mathfrak{g}(-1))$. Fix an $L$ orbit $\mathcal{O}$ on $\mathfrak{g}(-1)$. Just as above, we can define a micro-packet $\Pi^{g}_{F}(\mathcal{O})$ consisting of those $\pi_{F}(\gamma)$ such that the conormal bundle to $\mathcal{O}$ occurs in the characteristic cycle of $\textrm{per}(\gamma)$. Vogan has proposed a definition of Arthur packets as a certain special kind of micro-packets in the $p$-adic case. See [Vo] and [CuFK].
@@
For example, consider $G=\mathrm{GL}(n,\mathbb{C})$. In [BST], we found occurrences of the Kashiwara-Saito singularity in $K$ orbits on $\mathcal{P}$, and deduced the existence of reducible characteristic cycles. In the notation above, this gives examples of micro-packets of the form $\Pi^{g}_{\mathbb{R}}(Q)$ with more than one element for $\mathrm{GL}(n,\mathbb{R})$. From the proof of Theorem 3.1, one immediately deduces the occurrence of the Kashiwara-Saito singularity in the closure of an $L$ orbit on $\mathfrak{g}(-1)$, and the existence of micro-packets for $\mathrm{GL}(n,F)$ with more than one element, as discovered by [CuFK]. (It seems plausible that all of the interesting examples in [CuFK] are accounted for by matching with the real case as the choice of ordering on $\mathscr{P}$ varies.) In any event, since the theory in the real case is better developed, it is interesting to study other classical $p$-adic cases from this viewpoint. We would like to return to this elsewhere.
Finally, one especially interesting class of micro-packets for real groups are the special unipotent Arthur packets of [ABV, Chapter 27]. The construction of (6.1) applied to special unipotent packets for real classical groups should give rise to what one might call special unipotent packets for split $p$-adic classical groups. It would be interesting to compare this notion with the definition recently given by Ciubotaru, Mason-Brown, and Okada in [CMBO].

### v2

@@ Abstract.
Fix an integral semisimple element $\lambda$ in the Lie algebra $\mathfrak{g}$ of a complex reductive algebraic group $G$. Let $L$ denote the centralizer of $\lambda$ in $G$ and let $\mathfrak{g}(-1)$ denote the $-1$-eigenspace of $\mathop{\hbox{ad}}\nolimits(\lambda)$ in $\mathfrak{g}$. Under a natural hypothesis (which is always satisfied for classical groups), we embed the closure of each $L$ orbit on $\mathfrak{g}(-1)$ into the closure of an orbit of a symmetric subgroup $K$ containing $L$ on a partial flag variety for $G$. We use this to relate the local intersection homology of the latter orbit closures to the former orbit closures. This, in turn, relates multiplicity matrices for split real and $p$-adic groups. We also describe relationships between “microlocal packets” of representations of these groups.
@@
On the other hand, let $K$ denote the identity component of the fixed points in $G$ of the automorphism $\theta$ obtained by conjugation by $\exp(i\pi\lambda)$. Then $K$ acts with finitely many orbits on the partial flag variety $\mathcal{P}$ consisting of conjugates of the sum of the nonnegative eigenspaces of $\mathop{\hbox{ad}}\nolimits(\lambda)$. Once again, one can consider local intersection homology Poincaré polynomials. Vogan [Vo3] described a finite effective algorithm to compute them which has been implemented in the software package atlas. Evaluating these polynomials at 1 gives multiplicities of irreducibles in standard representation of the identity component of the split real form of the Langlands dual of $G$ [Vo4, ABV].
@@
Equation (1.1) generalizes the main geometric result for $\mathrm{GL}(n)$ of Ciubotaru-Trapa [CiT]; see Remark 3.3. In Remark 3.4, we also give analogous results equating the $p$-adic polynomials with Kazhdan-Lusztig polynomials arising from category $\mathcal{O}$, generalizing results of Zelevinsky [Z] for $\mathfrak{g}\mathfrak{l}(n)$ to all classical groups.
@@
Note also that we have assumed $K$ to be the identity component of the fixed points of $\theta$. This is perhaps unnatural from the point of view of representations of algebraic groups, and requires explanation. If we had instead worked with the potentially disconnected $K^{\prime}=G^{\theta}$, the orbits of $K^{\prime}$ on $\mathcal{P}$ can be reducible, and different irreducible components can contribute to the local intersection cohomology in ways that make the left-hand side of (1.1) a summand of the right-hand side. In any particular case, this is tractable to understand, but general statements in the presence of this kind of disconnectedness are somewhat cumbersome. See Remark 3.5.
@@
Let $G=\mathrm{GL}(n,{\mathbb{C}})$ and let $\mathfrak{t}$ denote the diagonal Cartan subalgebra. After a central shift, the integrality of $\lambda$ implies that we may assume $\lambda$ consists of integer entries. Then $K\simeq\mathrm{GL}(p,\mathbb{C})\times\mathrm{GL}(q,\mathbb{C})$ where $p$ is the number of even entries of $\lambda$ and $q$ is the number of odd entries. The symmetric pair $(G,K)$ corresponds to the real group $\mathrm{U}(p,q)$. This is the setting of [CiT].
@@
$\mathfrak{g}(-1)=\bigoplus_{i=1}^{\ell}\left[\mathfrak{l}_{i}\cap\overline{\mathfrak{u}}{}\right].$ (2.2)
@@
This is the map that is used in [CiT] to define $\mathcal{O}\mapsto Q_{\mathcal{O}}$. But note that we may take any ordering of the set $\mathscr{P}$ and thus define different maps $\mathcal{O}\mapsto Q_{\mathcal{O}}$ which will have the properties that we describe in Theorem 3.1. ∎
@@
Let $G$ be simple of type F4. Suppose $\lambda$ is one-half the middle element of a Jacobson-Morozov triple for the nilpotent orbit labeled in the Bala-Carter classification by F4(a3). The weighted Dynkin diagram of this orbit (in the standard Bourbaki order) is 0200. Thus $\mathfrak{p}{}$ is the maximal parabolic corresponding the long middle root on the Dynkin diagram of F4. The opposite nilradical $\overline{\mathfrak{u}}{}$ is a three-step nilpotent algebra whose first step is $\mathfrak{g}(-1)$. Since $\mathfrak{p}{}$ is maximal, the only possibility is for $\mathscr{P}$ to consist of a single algebra, namely all of $\mathfrak{g}$. But we have already remarked that $\mathfrak{g}\cap\overline{\mathfrak{u}}{}=\bar{\mathfrak{u}}{}$ properly contains $\mathfrak{g}(-1)$. Thus there is no choice of $\mathscr{P}$ that satisfies (2.2).∎
@@ 3.2. Statement of Main Results
Finally, fix $\mathcal{O}=L\cdot x$, $Q_{\mathcal{O}}=K\cdot\epsilon(x)$, and $\gamma\in\Xi(K,Y)$. The multiplicity of the conormal bundle to $Q_{\mathcal{O}}$ in the characteristic cycle of $\textrm{per}(\gamma)$ equals the multiplicity of the conormal bundle to $\mathcal{O}$ in $\textrm{per}(\varphi(\gamma))$,
$m^{mic}_{\mathcal{O}}(\textrm{per}(\varphi(\gamma)))=m^{mic}_{Q_{\mathcal{O}}}(\textrm{per}(\gamma)).$ (3.7)
@@
When $G=\mathrm{GL}(n,\mathbb{C})$ and the choice of $\mathscr{P}$ (and its ordering) is the one described in Example 2.5, then (3.6) in Theorem 3.1 is [CiT, Theorem 2.5]. ∎
@@
Note, however, that there are no nontrivial local systems in $\Xi(P,Y)$. So there is no hope in matching polynomials that arise for nontrivial local systems for $L$ orbits on $\mathfrak{g}(-1)$ with classical Kazhdan-Lusztig polynomials; cf. Remark 3.2. This is a reason for using $K$ orbits in order to match more general polynomials for $\mathfrak{g}(-1)$. ∎
@@
and that the analogous conclusion of (3.7) also holds. If we do not assume $Q_{\gamma}$ is irreducible, then (3.9) can fail; see [BT, Remark 3.6] for a discussion of an example in $\mathrm{SO}(8)$. The one place in the argument below that we need $Q_{\gamma}$ to be irreducible is to deduce the conclusion of Corollary 4.4.
@@ 4.3. Induced bundles
$m_{\mathcal{O}^{\prime}}(\textrm{per}(\gamma))=m_{\mathcal{O}}(\textrm{per}(\varphi_{2}(\gamma)).$ (4.14)
@@
In this section, we study a special class of examples that includes some cases (like Example 2.7) where there does not exist a set $\mathscr{P}$ satisfying (2.1) and (2.2). In these cases, we do not need to truncate the exponential map. Alternatively, one can think of this section as a kind of basic case, and the definition of $\epsilon$ in Definition 2.2 and some of the proofs of Section 4 as an induction using the bundle constructed in (4.5). This can be made precise, but isn’t necessary for our purposes here.
@@
$m^{mic}_{\mathcal{O}}(\textrm{per}(\varphi^{\prime}(\gamma)))=m^{mic}_{Q^{\prime}_{\mathcal{O}}}(\textrm{per}(\gamma)).$
@@
$[K\cap\bar{B}]\cdot\epsilon^{\prime}(x)$ is open and dense in $K\cdot\epsilon^{\prime}(x)$.
@@
To conclude, we sketch some consequences related to micro-packets of representations. Part of our motivation is to explain the appearance of the Kashiwara-Saito singularity discovered by Cunningham-Fiora-Kitt [CFK].
@@
Meanwhile, if we assume further that $\lambda$ is hyperbolic, then according to Lusztig’s classification of unipotent representation of graded affine Hecke algebras ([Lu1, Lu2, Lu3]), the set $\Xi(L,\mathfrak{g}(-1))$ parameterizes certain unipotent representations of the split $F=\mathbb{Q}_{p}$ form $G^{\prime}_{F}$ of the Langlands dual group $G^{\vee}$. (The subset $\Xi_{0}(L,\mathfrak{g}(-1))$ consisting of irreducible local systems of Springer type parametrizes irreducible unramified representations of $G^{\prime}_{F}$.) Write $\pi_{F}(\gamma)$ for the irreducible unipotent representation of $G^{\prime}_{F}$ corresponding to $\gamma\in\Xi(L,\mathfrak{g}(-1))$. Fix an $L$ orbit $\mathcal{O}$ on $\mathfrak{g}(-1)$. Just as above, we can define a micro-packet $\Pi^{g}_{F}(\mathcal{O})$ consisting of those $\pi_{F}(\gamma)$ such that the conormal bundle to $\mathcal{O}$ occurs in the characteristic cycle of $\textrm{per}(\gamma)$. Vogan has proposed a definition of Arthur packets as a certain special kind of micro-packets in the $p$-adic case. See [Vo] and [CFMMX].
@@
For example, consider $G=\mathrm{GL}(n,\mathbb{C})$. In [BST], we found occurrences of the Kashiwara-Saito singularity in $K$ orbits on $\mathcal{P}$, and deduced the existence of reducible characteristic cycles. In the notation above, this gives examples of micro-packets of the form $\Pi^{g}_{\mathbb{R}}(Q)$ with more than one element for $\mathrm{GL}(n,\mathbb{R})$. From the proof of Theorem 3.1, one immediately deduces the occurrence of the Kashiwara-Saito singularity in the closure of an $L$ orbit on $\mathfrak{g}(-1)$, and the existence of micro-packets for $\mathrm{GL}(n,F)$ with more than one element, as discovered by [CFK]. (It seems plausible that all of the interesting examples in [CFK] are accounted for by matching with the real case as the choice of ordering on $\mathscr{P}$ varies.) In any event, since the theory in the real case is better developed, it is interesting to study other classical $p$-adic cases from this viewpoint. We would like to return to this elsewhere.
Finally, one especially interesting class of micro-packets for real groups are the special unipotent Arthur packets of [ABV, Chapter 27]. The construction of (6.1) applied to special unipotent packets for real classical groups should give rise to what one might call special unipotent packets for split $p$-adic classical groups. It would be interesting to compare this notion with the definition recently given by Ciubotaru, Mason-Brown, and Okada in [CiMBO].